5. The Cobb–Douglas production function is

y " za

1 zb

2 a ' 0, b ' 0

Show that MP1 " ay/z1, MP2 " by/z2 .

What is the MRTS21? How does it vary with: (a) y; (b) z2 /z1?

Draw the isoquant m


1
Expert's answer
2021-07-22T10:53:34-0400

"Y=Z_1^{\alpha}Z_2^{\beta}\\
MP_1= \alpha Z_1^{\alpha -1}Z_2^{\beta}\\
MP_1= \frac{\alpha Z_1^{\alpha }Z_2^{\beta}}{Z_1}\\
MP_1= \frac{\alpha Y}{Z_1}\\"


"MP_2= \beta Z_1^{\alpha }Z_2^{\beta-1}\\
MP_2= \frac{\beta Z_1^{\alpha }Z_2^{\beta}}{Z_2}\\
MP_2= \frac{\beta Y}{Z_2}\\"


"MRTS= \frac{MP_1}{MP_2}\\
MRTS= \frac{\frac{\alpha Y}{Z_1}}{\frac{\beta Y}{Z_2}}\\
MRTS= \frac{\alpha}{\beta}*\frac{Z_2}{Z_1}\\"


a. MRTS would not be affected by the change in Y. As MRTS is not the function of Y.

b. As "\frac{Z_2}{Z_1}" changes MRTS changes.

An increase in "\frac{Z_2}{Z_1}" MRTS increases and vice versa.